OpenStudy (anonymous):

how to evaluate limit of a function with raised to power x as x tends to zero?

OpenStudy (kc_kennylau):

\[\LARGE\lim_{x\rightarrow0}f(x)^x=e^{\lim\limits_{x\rightarrow0}\ln f(x)^x}=e^{\lim\limits_{x\rightarrow0}x\ln f(x)}=\cdots\]